1.Suppose the monopolist faces the market demand function given by2144PQ.The AVCof the firm is given as AVC = Q½andthe firm has a fixed cost of $ 5a)determine equilibrium P&Qb)determine the maximumprofit

2.Given Qd = 60–15P + P², the (point) price elasticity of demand at a price of 5

3Primary and secondary schooling is free in state schools in our country. If parents aregiven a choice of schools fortheir children, there will be a shortage of places in popularschools. What methods could be used for dealing with this shortage? What are theirrelative merits?


1
Expert's answer
2021-05-03T10:48:24-0400

1. "Q = 144\/P^2" , "AVC = Q^{0.5}" , FC = $5.

a) In equilibrium MR = MC.

"P = 12\/Q^{0.5},"

"MR = TR'(Q) = 6\/Q^{0.5},"

"TC = 5 + Q^{1.5},"

"MC = TC'(Q) = 1.5Q^{0.5},"

"1.5Q^{0.5} = 6\/Q^{0.5}"

1.5Q = 6,

Q = 4 units,

P = 12/2 = 6.

b) The maximum profit is:

"TP = TR - TC = 6\u00d74 - (5 + 4^{1.5}) = 11."

2. Qd = 60 – 15P + P².

At a P = 5:

Qd = 60 - 15×5 + 25 = 10.

Q'd = -15 + 2P = -15 + 2×5 = -5.

The (point) price elasticity of demand is: "Ed = -5\u00d75\/10 = -2.5."

So, the demand is elastic.

3. The priority for children living close to such schools may be set or additional payment for children who live in other towns or cities.